OpenStudy (anonymous):

Write the quadratic equation having the given numbers as solutions -8 and -4. The quadratic equation is blank=0

OpenStudy (anonymous):

(x+8)(x+4)=0

OpenStudy (anonymous):

(x+8)(x+4)=0

OpenStudy (anonymous):

FOIL the factors

OpenStudy (anonymous):

\[x ^{2}+12x+32\]

OpenStudy (anonymous):

ok...need the same thing but with -3 and -1

OpenStudy (anonymous):

(x+3)(x+1)=0 FOIL

OpenStudy (anonymous):

\[x ^{2}+4x+3\]

OpenStudy (anonymous):

I thank you..I am going through the homework that I did and got a lot wrong and trying to figure out where i went wrong in all of it.

OpenStudy (anonymous):

=0

OpenStudy (anonymous):

ok